# Fast and Globally Consistent Normal Orientation based on the Winding Number Normal Consistency

### v1

@@ 3 Preliminaries
$F(y)$ $=$ $\int_{\partial\Omega}((\nabla\Phi)(y-x))\cdot n(x)\,\mathrm{d}S(x)$ (1)
$=$ $\left\{\begin{array}[]{lll}0&,&y\in\bar{\Omega}^{c}\\ 1/2&,&y\in\partial\Omega\\ 1&,&y\in\Omega\end{array}\right.$ (5)
@@
$\Phi(y)=\frac{1}{4\pi|y|},\quad\nabla\Phi(y)=-\frac{y}{4\pi|y|^{3}}.$ (6)
With a known normal vector field $n(x)$, the indicator function can be directly evaluated by Theorem 6. To numerically evaluate the winding number given a point set $\mathcal{P}=\{x_{j}\}_{j=1}^{N_{\mathcal{P}}}$ that samples the surface $\partial\Omega$, Lin et al. [31] adopt the following simple discretization:
$F(y)$ $=$ $\int_{\partial\Omega}((\nabla\Phi)(y-x))\cdot n(x)\,\mathrm{d}S(x)$ (7)
@@
$=$ $\sum_{j=1}^{N_{\mathcal{P}}}((\nabla\Phi)(y-x_{j}))\cdot\mu_{j}.$ (8)
Here, $\sigma(x_{j})$ denotes the local surface element area, $n(x_{j})$ denotes the unit vector and $\mu_{j}$ denotes their product, i.e., the oriented surface element at $x_{j}$. Lin et al. [31] further noticed that (5) in fact defines a linear equation in $\mu_{j}$:
$\sum_{j=1}^{N_{\mathcal{P}}}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}=1/2,\ \text{for all }x_{i}\in\mathcal{P}.$ (9)
Lin et al. [31] then rewrite Eq. (9) into a regularized symmetric system, which can then be efficiently solved with conjugate gradients [17]. Finally, $\mu_{j}$ are inserted back into Eq. (8) to compute the winding number field.
@@
Figure 3: The overall pipeline of our iterative algorithm. Each iteration consists of three steps: a gradient step w.r.t. a quadratic energy function, a normal update step based on the WNNC property and a rescaling step to ensure numerical stability. Our algorithm can generally produce an overall correct orientation at the first iteration, and gradually converges to details in a few dozens of iterations.
4 Method
4.1 Notations
@@
4.2 Winding Number Normal Consistency
In this section, we introduce our core contribution, the winding number normal consistency (WNNC). The WNNC property is based on a simple observation: If the given normals $\mu_{j}$ are consistently outward pointing, i.e., satisfying the assumptions in Theorem 6, then the winding number field $F(y;\mu)$ computed using these normals should be the indicator field of the corresponding shape. Consequently, the negative gradient of $F(y;\mu)$ at $x_{i}$ should point to the same direction as $\mu_{i}$. Fig. 2 illustrates two cases, where one starts with (possibly incorrectly oriented) normals $\mu$, evaluates $F(y;\mu)$ and then resamples its negative gradient at each $x_{i}$. Note that if the original normals $\mu$ are correctly oriented, the resampled ones would still be aligned with them. The WNNC property can be expressed mathematically as follows: If $\mu_{j}$ are consistently outward pointing, then
$\mu_{i}=\lambda_{i}(-\nabla F(x_{i};\mu))\quad\textrm{for some}\quad\lambda_{i}>0.$ (10)
@@ 4.2 Winding Number Normal Consistency
$\nabla F(y;\mu)$ $=$ $\nabla\Big{(}\sum_{j}((\nabla\Phi)(y-x_{j}))\cdot\mu_{j}\Big{)}$ (11)
$=$ $\sum_{j}((H\Phi)(y-x_{j}))\cdot\mu_{j},$
@@
$H\Phi(y)=-\frac{I_{3}}{4\pi|y|^{3}}+\frac{3}{4\pi|y|^{5}}\left[\begin{matrix}y_{1}\\ y_{2}\\ y_{3}\end{matrix}\right][y_{1}\ y_{2}\ y_{3}].$ (12)
@@
Finally we can conclude the WNNC property in a simple mathematical form: If $\mu_{j}$ are consistently outward pointing, then
$\mu_{i}=\lambda_{i}(-\nabla F(y_{i};\mu))=\lambda_{i}G(\mu)_{i}\quad\textrm{for some}\quad\lambda_{i}>0.$ (13)
4.3 WNNC-based Iterative Algorithm
@@
In this section, we describe our algorithm that utilizes the WNNC property Eq. (13) to achieve fast and high-quality convergence. At first sight, one may think it is natural to define an objective function to minimize the discrepancy between the directions of $\mu$ and $G(\mu)$, e.g., using the cosine similarity between them. However, this would lead to a non-linear problem that may be complicated to solve. We observe that the WNNC condition Eq. (13) seems to hint $\mu$ is the fixed point of the mapping $\mu\mapsto G(\mu)$ (up to per-point scaling). Inspired by fixed-point iteration algorithm in classical numerical analysis, we attempt to adopt an iterative scheme for solving normals using the WNNC property, where we update $\mu$ using $\mu\mapsto G(\mu)$. We call $\mu\mapsto G(\mu)$ a WNNC update.
@@
Despite the similarity between WNNC updates and fixed-point iterations, merely applying the fixed-point iteration algorithm is infeasible due to the following reasons. (a) A trivial solution $\mu=0$ is clearly a fixed point of $\mu\mapsto G(\mu)$. (b) Even if we avoid $\mu=0$, there is no guarantee of convergence when using WNNC updates. In practice we also find the WNNC update alone is insufficient for convergence. (c) The WNNC property involves only the normal directions, but says nothing about the scale changes $\lambda_{i}$. Thus, $\mu$ is not a fixed point of $\mu\mapsto G(\mu)$ in a traditional sense. Furthermore, we empirically found $\lambda_{i}$ to range from $10^{4}$ to $10^{10}$ for typical datasets, which would blow up numerically in only a few iterations.
@@
Since WNNC updates alone are insufficient, we propose to incorporate the constraints Eq. (9) from PGR [31] as an additional objective function:
$E=\sum_{i}\Big{|}\sum_{j=1}^{N_{\mathcal{P}}}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}-1/2\Big{|}^{2},$ (14)
@@
Again, to simplify notations, we define the mapping $\mu\mapsto s$, where $s_{i}:=F(x_{i};\mu)$, as $s=A(\mu)$. In other words, $\mu\mapsto s=A(\mu)$ evaluates the winding number field at each $x_{i}$ with given normals $\mu$. In this way, we can rewrite $E$ as
$E(\mu;A,b)=\|A(\mu)-b\|^{2},\ \textrm{with}\ b=(1/2,\cdots,1/2)^{T}.$ (15)
@@
Our iterative algorithm aims to update normals using both WNNC updates $\mu\mapsto G(\mu)$ and $E(\mu;A,b)$. More specifically, our algorithm starts with an all-zero initial vector field $\mu=0$. During each iteration, we first take a gradient step w.r.t. $E(\mu;A,b)$, and then update the normals using an WNNC update. To solve the scale change problem during WNNC updates mentioned above, we rescale each $\mu_{i}$ to have the same length after each WNNC update. The algorithm is summarized as Algorithm 1, and Fig. 3 presents a graphical overview of our iterative pipeline.
@@
4.4 Resolving Singularities
The steps in Algorithm 1 are none other than evaluating expressions of the form $(\nabla\Phi)(x_{i}-x_{j})$ and $(H\Phi)(x_{i}-x_{j})$. However, both expressions become singular when $x_{i}$ approaches $x_{j}$. Using their raw forms inevitably brings numerical instability. Following previous work [32], we set their values to 00 if $\|x_{i}-x_{j}\|<w$, where $w>0$ is a constant called the smoothing width. This change affects the operators $A$ and $G$. We denote their modified versions by $A_{w}$ and $G_{w}$, respectively. As reported in previous work [32, 31], a small $w$ retains more details but is less robust to noise and other imperfections, while a large $w$ leads to an overly smooth winding number field. In practice, we find that using a fixed width as in PGR [31] either leads to oversmoothing or cannot ensure global consistency. We thus propose a width scheduling scheme, where we start with a large $w=w_{2}$ and linearly decreases it to a small $w=w_{1}$ as the algorithms proceeds. In practice, This allows the iterative algorithm to remain stable during early iterations, and to enhance details in late iterations. The final algorithm is described in Algorithm 3.
@@ 4.4 Resolving Singularities
Figure 4: Illustration of the treecode-based approximation algorithm [4]. (a) We wish to sum up interactions between multiple source points and the target point. (b) The space is partitioned by an octree hosting the source points, and an representative point is computed in each node (not shown here). (c) Based on the proximity between the target point and the octree nodes, representative points are chosen. Attributes of the source points are aggregated to their corresponding representative points. (d) For far-field interactions, representative points are used instead of the original source points.
4.5 Treecode-based Acceleration on GPU
As described in Algorithm 2 and Algorithm 3, our iterative scheme mainly involves evaluating the three linear operators $\mu\mapsto A(\mu)$, $p\mapsto A^{T}(p)$ and $\mu\mapsto G(\mu)$. A naive method for evaluating each of these operator has time complexity $O(N_{\mathcal{P}}^{2})$. The computational time becomes prohibitively long as the number of input points grows.
@@ 4.5 Treecode-based Acceleration on GPU
To build representative sources, we first build an octree $\mathcal{T}$ to spatially partition the input point set. The partitioning stops if the each leaf node contains only one point or if the user-specified maximum depth $D$ is reached. Then, at each iteration, each octree node accumulates point attributes (such as $\mu_{i}$) from the points it contains and select a representative point as a weighted average of all points in the node. The representative point of each node is determined as follows. Let $B$ be some node of the octree and let $\mathcal{P}_{B}$ be the subset of input points that lie in $B$. Let $\nu_{j}$ be an attribute vector of $x_{j}$ that we wish to propagate to $x_{i}$ through an interaction (e.g., $\nu_{j}=\mu_{j}$ when evaluating $\mu\mapsto A(\mu)$ or $\mu\mapsto G(\mu)$, and $\nu_{j}=s_{j}$ when evaluating $s\mapsto A^{T}s$). The $\nu$-representative of the node $B$ is located at
$x_{B,\nu}=\sum_{x_{i}\in\mathcal{P}_{B}}\frac{|\nu_{i}|}{\sum_{x_{j}\in\mathcal{P}_{B}}|\nu_{j}|}x_{i},$ (16)
@@
$\sum_{x_{j}\ {\rm inside}\ B}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}.$ (18)
@@
To get $s_{i}$ in $s=A(\mu)$, simply call ${\rm apply\_A}(x_{i},\mathcal{T}_{\rm root})$. Note that $c$ and $w$ are hyperparameters. $c$ determines whether to use the representative as an approximation, and $w$ is the smoothing width in Sec. 4.4. Finally, we parallelize all evaluatations ${\rm apply\_A}(x_{i},\mathcal{T}_{\rm root})$ across all $x_{i}$. Other operators $A^{T}$ and $G$ are accelerated in the same way.
@@
We remark that the idea of using a tree structure to approximate far-field interactions dates back to Barnes and Hut [4], which was originally proposed to accelerate the gravitational $N$-body problem. The tree-based algorithm is now referred to as treecode [45]. The fast multipole method (FMM) [14] pushes this idea further by using representatives for target points as well and by using expansions for propagation kernels for precision control. Previous works [3, 32] have also utilized these acceleration techniques to accelerate evaluating winding numbers. While treecodes and FMM are well-studied with publicly available libraries, to the best of our knowledge, there is currently no open-source implementation of treecode or FMM that suits our requirements: (a) GPU-based; (b) support for the operators used in our algorithm: $A,A^{T}$ and $G$. For this reason, we implement our own CUDA kernels for our algorithm.
5 Experiments
5.1 Experiment Setup
5.1.1 Implementation Details

### v2

@@ 3. Preliminaries
(1) $F(y)$ $=$ $\int_{\partial\Omega}((\nabla\Phi)(y-x))\cdot n(x)\,\mathrm{d}S(x)$
(5) $=$ $\left\{\begin{array}[]{lll}0&,&y\in\bar{\Omega}^{c}\\ 1/2&,&y\in\partial\Omega\\ 1&,&y\in\Omega\end{array}\right.$
@@
(6) $\Phi(y)=\frac{1}{4\pi|y|},\quad\nabla\Phi(y)=-\frac{y}{4\pi|y|^{3}}.$
With a known normal vector field $n(x)$, the indicator function can be directly evaluated by Theorem 6. To numerically evaluate the winding number given a point set $\mathcal{P}=\{x_{j}\}_{j=1}^{N_{\mathcal{P}}}$ that samples the surface $\partial\Omega$, Lin et al. (2022) adopt the following simple discretization:
(7) $F(y)$ $=$ $\int_{\partial\Omega}((\nabla\Phi)(y-x))\cdot n(x)\,\mathrm{d}S(x)$
@@
(8) $=$ $\sum_{j=1}^{N_{\mathcal{P}}}((\nabla\Phi)(y-x_{j}))\cdot\mu_{j}.$
Here, $\sigma(x_{j})$ denotes the local surface element area, $n(x_{j})$ denotes the unit vector and $\mu_{j}$ denotes their product, i.e., the oriented surface element at $x_{j}$. Lin et al. (2022) further noticed that (5) in fact defines a linear equation in $\mu_{j}$:
(9) $\sum_{j=1}^{N_{\mathcal{P}}}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}=1/2,\ \text{for all }x_{i}\in\mathcal{P}.$
Lin et al. (2022) then rewrite Eq. (9) into a regularized symmetric system, which can then be efficiently solved with conjugate gradients (Hestenes and Stiefel, 1952). Finally, the $\mu_{j}$’s are inserted back into Eq. (8) to compute the winding number field.
@@
Figure 3. The overall pipeline of our iterative algorithm. Each iteration consists of three steps: a gradient step w.r.t. a quadratic energy function proposed in Lin et al. (2022), a normal update step based on the WNNC property, and a rescaling step to ensure numerical stability. Our algorithm can generally produce an overall correct orientation at the first iteration, and gradually converges to details in a few dozens of iterations.
[…]
Illustration of the pipeline of our iterative algorithm.
4. Method
4.1. Notations
@@
4.2. Winding Number Normal Consistency
In this section, we introduce our core contribution, the winding number normal consistency (WNNC). The WNNC property is based on a simple observation: If the given normals $\mu_{j}$ are consistently outward-pointing, i.e., satisfying the assumptions in Theorem 6, then the winding number field $F(y;\mu)$ computed using these normals should be the indicator field of the corresponding shape. Consequently, the negative gradient of $F(y;\mu)$ at $x_{i}$ should point to the same direction as $\mu_{i}$. Fig. 2 illustrates two cases, where one starts with (possibly incorrectly oriented) normals $\mu$, evaluates $F(y;\mu)$ and then resamples its negative gradient at each $x_{i}$. Note that if the original normals $\mu$ are correctly oriented, the resampled ones would still be aligned with them. The WNNC property can be expressed mathematically as follows: If $\mu_{j}$ are consistently outward-pointing, then
(10) $\mu_{i}=\lambda_{i}(-\nabla F(x_{i};\mu))\quad\textrm{for some}\quad\lambda_{i}>0.$
@@ 4.2. Winding Number Normal Consistency
(11) $\nabla F(y;\mu)=\nabla\Big{(}\sum_{j}((\nabla\Phi)(y-x_{j}))\cdot\mu_{j}\Big{)}=\sum_{j}((H\Phi)(y-x_{j}))\cdot\mu_{j},$
@@
(12) $H\Phi(y)=-\frac{I_{3}}{4\pi|y|^{3}}+\frac{3}{4\pi|y|^{5}}\left[\begin{matrix}y_{1}\\ y_{2}\\ y_{3}\end{matrix}\right][y_{1}\ y_{2}\ y_{3}].$
@@
Finally, we can conclude the WNNC property in a simple mathematical form: If $\mu_{j}$ are consistently outward-pointing, then
(13) $\mu_{i}=\lambda_{i}(-\nabla F(y_{i};\mu))=\lambda_{i}G(\mu)_{i}\quad\textrm{for some}\quad\lambda_{i}>0.$
4.3. WNNC-based Iterative Algorithm
@@
In this section, we describe our algorithm that utilizes the WNNC property Eq. (13) to achieve fast and high-quality convergence. At first sight, one may think it is natural to define an objective function to minimize the discrepancy between the directions of $\mu$ and $G(\mu)$, e.g., using the cosine similarity between them. However, this would lead to a non-linear problem that may be complicated to solve. We observe that the WNNC condition Eq. (13) seems to hint $\mu$ is the fixed point of the mapping $\mu\mapsto G(\mu)$ (up to per-point scaling). Inspired by fixed-point iteration algorithms in classical numerical analysis, we attempt to adopt an iterative scheme for solving normals using the WNNC property, where we update $\mu$ using $\mu\mapsto G(\mu)$. We call $\mu\mapsto G(\mu)$ a WNNC update.
@@
Despite the similarity between WNNC updates and fixed-point iterations, merely applying the fixed-point iteration algorithm is infeasible due to the following reasons. (a) A trivial solution $\mu=0$ is clearly a fixed point of $\mu\mapsto G(\mu)$. (b) Both consistently outward-pointing normals and consistently inward-pointing normals are invariant under WNNC updates (up to scaling), which suggests the WNNC alone may not be able to flip incorrect normals. In practice, we also find merely using WNNC updates is insufficient for convergence (see Sec. 6.1.3). (c) The WNNC property involves only the normal directions, but says nothing about the scale changes $\lambda_{i}$. Thus, $\mu$ is not a fixed point of $\mu\mapsto G(\mu)$ in a traditional sense. Furthermore, we empirically found $\lambda_{i}$ to range from $10^{4}$ to $10^{10}$ for typical datasets, which would blow up numerically in only a few iterations.
@@
Since WNNC updates alone are insufficient, we propose to incorporate the constraints Eq. (9) from PGR (Lin et al., 2022) as an additional objective function:
(14) $E=\sum_{i}\Big{|}\sum_{j=1}^{N_{\mathcal{P}}}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}-1/2\Big{|}^{2},$
@@
We remark that, as discussed in Lin et al. (2022), the above equation can resolve the inside/outside ambiguity and encourages the resulting normals to be consistently outward-pointing. Again, to simplify notations, we define the mapping $\mu\mapsto s$, where $s_{i}:=F(x_{i};\mu)$, as $s=A(\mu)$. In other words, $\mu\mapsto s=A(\mu)$ evaluates the winding number field at each $x_{i}$ with given normals $\mu$. In this way, we can rewrite $E$ as
(15) $E=E(\mu;A,b)=\|A(\mu)-b\|^{2},\quad\textrm{where}\quad b=(1/2,\cdots,1/2)^{T}.$
@@
Our iterative algorithm aims to update normals using both WNNC updates $\mu\mapsto G(\mu)$ and $E(\mu;A,b)$. More specifically, our algorithm starts with an all-zero initial vector field $\mu=0$. During each iteration, we first take a gradient step w.r.t. $E(\mu;A,b)$, and then update the normals using an WNNC update. To solve the scale change problem during WNNC updates mentioned above, we rescale each $\mu_{i}$ to have the same length as before each WNNC update. The algorithm is summarized as Algorithm 1, and Fig. 3 presents a graphical overview of our iterative pipeline.
@@
4.4. Resolving Singularities
The steps in Algorithm 1 are none other than evaluating expressions of the form $(\nabla\Phi)(x_{i}-x_{j})$ and $(H\Phi)(x_{i}-x_{j})$. However, both expressions become singular when $x_{i}$ approaches $x_{j}$. Using their raw forms inevitably brings numerical instability. Following previous work (Lu et al., 2018), we set their values to 00 if $\|x_{i}-x_{j}\|<w$, where $w>0$ is a constant called the smoothing width. This change affects the operators $A$ and $G$. We denote their modified versions by $A_{w}$ and $G_{w}$, respectively. As reported in previous work (Lu et al., 2018; Lin et al., 2022), a small $w$ retains more details but is less robust to noise and other imperfections, while a large $w$ leads to an overly smooth winding number field. In practice, we find that using a fixed width as in PGR (Lin et al., 2022) either leads to oversmoothing or cannot ensure global consistency. We thus propose a width scheduling scheme, where we start with a large $w=w_{2}$ and linearly decreases it to a small $w=w_{1}$ as the algorithm proceeds. In practice, this allows the iterative algorithm to remain stable during early iterations, and to enhance details in late iterations. The final algorithm is described in Algorithm 3.
@@ 4.4. Resolving Singularities
Figure 4. Illustration of the treecode-based approximation algorithm (Barnes and Hut, 1986). (a) We wish to sum up interactions between multiple source points and the target point. (b) The space is partitioned by an octree hosting the source points, and a representative point is computed in each node (not shown here). (c) Based on the proximity between the target point and the octree nodes, representative points are chosen. Attributes of the source points are aggregated to their corresponding representative points. (d) For far-field interactions, representative points are used instead of the original source points.
4.5. Treecode-based Acceleration on GPU
As described in Algorithm 2 and Algorithm 3, our iterative scheme mainly involves evaluating the three linear operators $\mu\mapsto A(\mu)$, $p\mapsto A^{T}(p)$ and $\mu\mapsto G(\mu)$. A naive method for evaluating each of these operators has time complexity $O(N_{\mathcal{P}}^{2})$. The computational time becomes prohibitively long as the number of input points grows.
@@ 4.5. Treecode-based Acceleration on GPU
To build representative sources, we first build an octree $\mathcal{T}$ to spatially partition the input point set. The partitioning stops if the node contains only one point or if the user-specified maximum depth $D$ is reached. Then, at each iteration, each octree node accumulates point attributes (such as $\mu_{i}$) from the points it contains and select a representative point as a weighted average of all points in the node. The representative point of each node is determined as follows. Let $B$ be some node of the octree and let $\mathcal{P}_{B}$ be the subset of input points that lie in $B$. Let $\nu_{j}$ be an attribute vector of $x_{j}$ that we wish to propagate to $x_{i}$ through an interaction (e.g., $\nu_{j}=\mu_{j}$ when evaluating $\mu\mapsto A(\mu)$ or $\mu\mapsto G(\mu)$, and $\nu_{j}=s_{j}$ when evaluating $s\mapsto A^{T}s$). The $\nu$-representative of the node $B$ is located at
(16) $x_{B,\nu}=\sum_{x_{i}\in\mathcal{P}_{B}}\frac{|\nu_{i}|}{\sum_{x_{j}\in\mathcal{P}_{B}}|\nu_{j}|}x_{i},$
@@
(18) $\sum_{x_{j}\ {\rm inside}\ B}(\nabla\Phi)(x_{i}-x_{j})\cdot\mu_{j}.$
@@
To get $s_{i}$ in $s=A(\mu)$, simply call ${\rm apply\_A}(x_{i},\mathcal{T}_{\rm root})$. Note that $c$ and $w$ are hyperparameters. $c$ determines whether to use the representative as an approximation, and $w$ is the smoothing width in Sec. 4.4. Finally, we parallelize all evaluations ${\rm apply\_A}(x_{i},\mathcal{T}_{\rm root})$ across all $x_{i}$. Other operators $A^{T}$ and $G$ are accelerated in the same way.
@@
We remark that the idea of using a tree structure to approximate far-field interactions dates back to Barnes and Hut (1986), which was originally proposed to accelerate the gravitational $N$-body problem. The tree-based algorithm is now referred to as treecode (Yokota and Barba, 2011). The fast multipole method (FMM) (Greengard and Rokhlin, 1997) pushes this idea further by using representatives for target points as well and by using expansions for propagation kernels for precision control. Previous works (Barill et al., 2018; Lu et al., 2018) have also utilized these acceleration techniques to accelerate evaluating winding numbers. While treecodes and FMM are well-studied with publicly available libraries, to the best of our knowledge, there is currently no open-source implementation of treecode or FMM that suits our requirements: (a) GPU-based; (b) support for the operators used in our algorithm: $A,A^{T}$ and $G$. For this reason, we implement our own CUDA kernels and integrate them in PyTorch (Paszke et al., 2019; Ansel et al., 2024). Note that the $A^{T}$ operator is exactly the backward function for $A$, which means our code also supports differentiable programming involving winding numbers. We hope our implementation will further advance the research into winding numbers.
5. Experiments
5.1. Experiment Setup
5.1.1. Implementation Details